# On the Identifying Power of Generalized Monotonicity

### v2

@@ 3.1 Identified Sets for $\mathbf{Q}\subseteq\mathbf{Q}_{1,2}^{\ast}$
Suppose $\mathbf{Q}\subseteq\mathbf{Q}_{1,2}^{\ast}$ and $\mathbf{Q}|_{Y}$ equals the entire set of distributions on $\mathcal{Y}^{|\mathcal{D}|}$. Further suppose $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$. Then,
$\Theta_{0}(P,\mathbf{Q})=\prod_{d\in\mathcal{D}}\left[\beta_{d|z^{*}(d)}+y^{L}(1-\sum_{y\in\mathcal{Y}}p_{yd|z^{*}(d)})),~{}\beta_{d|z^{*}(d)}+y^{U}(1-\sum_{y\in\mathcal{Y}}p_{yd|z^{*}(d)}))\right]~{}.$ (5)
@@
Under the instrument exogeneity and monotonicity assumptions of Imbens and Angrist (1994), Balke and Pearl (1993, 1997) found the same form of the identified set for $\theta(Q)$ as (5) when $\mathcal{Y}=\mathcal{D}=\mathcal{Z}=\{0,1\}$. Theorem 3.2 therefore generalizes the result of Balke and Pearl (1993, 1997) to more than two treatment arms and instrument values, to outcomes taking more than two values, and to show that the same identified set holds when imposing possibly stronger restrictions on potential treatments than generalized monotonicity.
@@
$\biggl{[}(\beta_{j|z^{*}(j)}-\beta_{k|z^{*}(k)})+(y^{L}-y^{U})+y^{U}\sum_{y\in\mathcal{Y}}p_{yk|z^{*}(k)}-y^{L}\sum_{y\in\mathcal{Y}}p_{yj|z^{*}(j)}~{},\\ ~{}(\beta_{j|z^{*}(j)}-\beta_{k|z^{*}(k)})+(y^{U}-y^{L})+y^{L}\sum_{y\in\mathcal{Y}}p_{yk|z^{*}(k)}-y^{U}\sum_{y\in\mathcal{Y}}p_{yj|z^{*}(j)}\biggr{]}~{}.$ (6)
@@
Therefore, when imposing generalized monotonicity, as well as when imposing any restriction implying generalized monotonity, the ATE of $j$ versus $k$ is only identified “at infinity” (Heckman, 1990; Andrews and Schafgans, 1998) in the sense that identification requires
$\min_{z\in\mathcal{Z}}P\{D\neq j\mid Z=z\}=\min_{z\in\mathcal{Z}}P\{D\neq k\mid Z=z\}=0~{}.$
In other words, identification of the the ATE of $j$ versus $k$ under generalized monotonicity or under any restriction implying generalized monotonicity requires that there is some value of the instrument such that everyone takes treatment $j$ at that value of the instrument, and some value of the instrument such that everyone takes treatment $k$ at that value of the instrument.
@@ 3.2 Identifying Power of Generalized Monotonicity
Theorem 3.2 above establishes that, for possibly multi-valued $Y$, $D$, and $Z$, and any model $\mathbf{Q}\subseteq\mathbf{Q}_{1,2}^{\ast}$ such that $\mathbf{Q}|_{Y}$ equals the entire set of distributions on $\mathcal{Y}^{|\mathcal{D}|}$, if $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$, then $\Theta_{0}(P,\mathbf{Q})$ equals (5). We now show that $\Theta_{0}(P,\mathbf{Q}_{1,2}^{\ast})=\Theta_{0}(P,\mathbf{Q}_{1}^{\ast})$ as long as $\mathbf{Q}_{0}(P,\mathbf{Q}_{1,2}^{\ast})\neq\emptyset$, so that the identified set for $\theta(Q)$ assuming instrument exogeneity and generalized monotonicity coincides with the identified set assuming instrument exogeneity alone, as long as both assumptions are consistent with the distribution of the observed data. Our result therefore generalizes Balke and Pearl (1993, 1997), which study the case of binary $Y$, $Z$, and $D$.
@@
Note Assumption 3.1 is weaker than instrument exogeneity in Assumption 2.1, and does not imply (2). Let $\mathbf{Q}_{3}^{\ast}$ denote the set of all $Q$ that satisfies Assumption 3.1. Following Robins (1989) and Manski (1990), the following lemma derives the identified set for $\theta(Q)$ under mean independence:
@@
Suppose $\mathbf{Q}_{0}(P,\mathbf{Q}_{3}^{\ast})\neq\emptyset$. Then,
$\Theta_{0}(P,\mathbf{Q}_{3}^{\ast})=\prod_{d\in\mathcal{D}}\left[\max_{z\in\mathcal{Z}}\{\beta_{d|z}+y^{L}(1-\sum_{y\in\mathcal{Y}}p_{yd\mid z})\},~{}\min_{z\in\mathcal{Z}}\{\beta_{d|z}+y^{U}(1-\sum_{y\in\mathcal{Y}}p_{yd\mid z})\}\right].$ (7)
The following lemma, which relies on the observation in Lemma 2.1, establishes the equivalence between the identified sets in (5) and (7) when $\mathbf{Q}_{0}(P,\mathbf{Q}_{1,2}^{\ast})\neq\emptyset$.
@@
Suppose $\mathbf{Q}_{0}(P,\mathbf{Q}_{1,2}^{\ast})\neq\emptyset$. Then, the sets in (5) and (7) coincide.
@@
Suppose $\mathbf{Q}_{0}(P,\mathbf{Q}_{1,2}^{\ast})\neq\emptyset$. Then, $\Theta_{0}(P,\mathbf{Q}_{1,2}^{\ast})=\Theta_{0}(P,\mathbf{Q}_{1}^{\ast})$.
@@
We now consider some restrictions on potential treatments that have been considered previously in the literature. In each case, we show $\mathbf{Q}\subseteq\mathbf{Q}_{1,2}^{\ast}$; in particular, these restrictions satisfy generalized monotonicity. We emphasize that frequently $\mathbf{Q}\subsetneqq\mathbf{Q}_{1,2}^{\ast}$. In what follows, it is implicitly understood that $\mathbf{Q}|_{Y}$ equals the entire set of distributions on $\mathcal{Y}^{|\mathcal{D}|}$, so that the model imposes no restriction on potential outcomes. Thus, in each example Theorem 3.1 applies and such restrictions do not provide any identifying power for average potential outcomes or ATEs.
@@
Consider a multi-arm randomized controlled trial (RCT) with noncompliance, where $Z=d$ denotes random assignment to treatment $d$, $D_{d}=d$ denotes that the subject would comply with assignment if assigned to treatment $d$, and $|\mathcal{D}|=|\mathcal{Z}|$. More generally, not necessarily in the context of an RCT, one can interpret $Z=d$ as encouragement to treatment $d$ and interpret $D_{d}=d$ as the subject would take treatment $d$ if encouraged to do so. In this example, $Q$ satisfies Assumption 2.1 because $Z$ is randomly assigned. We may generalize the “no-defier” restriction of of Angrist et al. (1996) as follows:
@@
i.e., there is zero probability that a subject would not take treatment $d$ if assigned to (encouraged to take) $d$ but would take $d$ if assigned (encouraged) to some other treatment $d^{\prime}\neq d$. If $Q$ satisfies this generalized no-defier restriction, then Assumption 2.2 holds with $z^{*}(d)=d$ for all $d$. This no-defier restriction in particular holds in the context of an RCT with “one-sided non-compliance,” where we assume
@@
$Q\{D_{1}=2\mid D_{0}\neq D_{1}\}=1~{}.$ (8)
The condition in (8) states that if the choice of a family changes upon receiving a Head Start offer, then they must choose Head Start when receiving the offer. In other words, it cannot be the case that upon receiving a Head Start offer, a family switches from no preschool to preschools other than Head Start, or the other way around. Assumption 2.2 then holds with $z^{*}(0)=z^{*}(1)=0$ and $z^{*}(2)=1$. To see this, note that (8) implies $Q\{D_{0}\neq D_{1},D_{1}\neq 2\}=0$ and thus
$Q\{D_{0}\neq 0,D_{1}=0\}$ $=Q\{D_{0}\neq D_{1},D_{1}=0\}=0~{},$
$Q\{D_{0}\neq 1,D_{1}=1\}$ $=Q\{D_{0}\neq D_{1},D_{1}=1\}=0~{},$
$Q\{D_{1}\neq 2,D_{0}=2\}$ $\leq Q\{D_{0}\neq D_{1},D_{1}\neq 2\}=0~{}.$
Note in this example Assumption 2.2 still holds although $|\mathcal{Z}|<|\mathcal{D}|$.
Example 4.4.
@@
Assumption 2.2 holds for any $Q$ that satisfies (9). To see this, note that unordered monotonicity immediately implies that, for each $d\in\mathcal{D}$, $\mathbbm{1}\{D_{z}=d\}$ is weakly ordered across all $z\in\mathcal{Z}$ (w.p.1), and Assumption 2.2 is immediately seen to hold by taking $z^{\ast}(d)$ to be the maximum with respect to this order. Note, however, that $\mathcal{Z}^{*}(d)$ may not be a singleton unless some inequalities in (9) are strict.
@@
Although unordered monotonicity implies Assumption 2.2, the converse is generally false. For example, suppose $\mathcal{Z}=\{0,1,2,3\}$ and $\mathcal{D}=\{0,1\}$. Suppose $\mathbbm{1}\{D_{3}=1\}\geq\mathbbm{1}\{D_{z}=1\}$ w.p.1 under $Q$ for $z\neq 3$ and $\mathbbm{1}\{D_{0}=0\}\geq\mathbbm{1}\{D_{z}=0\}$ w.p.1 under $Q$ for $z\neq 0$, but $Q\{D_{1}=1,D_{2}=0\}$ and $Q\{D_{1}=0,D_{2}=1\}$ are both strictly positive. Then Assumption 2.2 holds with $z^{*}(0)=0$ and $z^{*}(1)=3$, but unordered monotonicity fails. In particular, $\mathbbm{1}\{D_{1}=1\}$ and $\mathbbm{1}\{D_{2}=1\}$ are not ordered, thus violating (9). We thus conclude that if $\mathbf{Q}$ is defined as the set of distributions that satisfy instrument exogeneity and unordered monotonicity, then $\mathbf{Q}\subsetneqq\mathbf{Q}_{1,2}^{\ast}$.
@@
for $g:\mathcal{Z}\times\mathcal{D}\to\Re$ where $\Re$ is the set of real numbers and a random vector $(U_{d}:d\in\mathcal{D})$, whose distribution is absolutely continuous with respect to the Lebesgue measure on $\Re^{|\mathcal{D}|}$ and $Z\perp\!\!\!\perp((U_{d}:d\in\mathcal{D}),(Y_{d}:d\in\mathcal{D}))$. Hence, $Q$ satisfies Assumption 2.1 by construction. Let $\mathbf{Q}$ denote the set of distributions that are consistent with $(D_{z}:z\in\mathcal{Z})$ being determined by (10) for some $g$ and $(U_{d}:d\in\mathcal{D})$ satisfying these requirements. The model $\mathbf{Q}$ is called an additive random utility model (ARUM). A sufficient condition for $Q\in\mathbf{Q}$ to satisfy Assumption 2.2 is that for each $d\in\mathcal{D}$ there exists $z^{*}(d)\in\mathcal{Z}$ such that
$g(z^{*}(d),d)-g(z^{*}(d),d^{\prime})>g(z,d)-g(z,d^{\prime})\text{ for all }d^{\prime}\neq d\text{ and }z\neq z^{*}(d)~{}.$ (11)
We refer to the requirement in (11) as uniform targeting of treatment $d$. The terminology is intended to reflect that there is a value of the instrument that maximizes the gains (in terms of $g$) of choosing treatment $d$ versus any other treatment $d^{\prime}$ uniformly across these other possible values of the treatment. In this sense, that value of the instrument targets treatment $d$ uniformly. We now argue by contradiction that (11) implies (3); hence, if (11) holds for all $d\in\mathcal{D}$, then $Q$ satisfies Assumption 2.2. To this end, suppose that, with positive probability, $D_{z^{*}(d)}=d^{\prime}\neq d$ but $D_{z^{\prime}}=d$ for $z^{\prime}\neq z^{*}(d)$. Then,
$g(z^{*}(d),d^{\prime})+U_{d^{\prime}}$ $\geq g(z^{*}(d),d)+U_{d}~{},$
@@
$g(z^{*}(d),d)-g(z^{*}(d),d^{\prime})\leq g(z^{\prime},d)-g(z^{\prime},d^{\prime})~{},$
which violates (11). A particular example when $|\mathcal{Z}|\geq|\mathcal{D}|$ that satisfies (11) with $z^{*}(d)=d$ after a suitable relabelling is
@@
Example 4.6.
$D_{z}=\mathbbm{1}\{g_{10}(z)+U_{10}\geq 0\}~{}.$ (13)
Let $\overline{\mathcal{Z}}=\operatorname*{argmax}_{z\in\mathcal{Z}}\{g_{10}(z)\}$, and let $\underline{\mathcal{Z}}=\operatorname*{argmin}_{z\in\mathcal{Z}}\{g_{10}(z)\}$. Then, $Q$ satisfies Assumption 2.2 with $\mathcal{Z}^{*}(1)=\overline{\mathcal{Z}}$ and $\mathcal{Z}^{*}(0)=\underline{\mathcal{Z}}$. To contrast with Example 4.5, note that (11) holds if and only if $\overline{\mathcal{Z}}$ and $\underline{\mathcal{Z}}$ are both singletons.
@@
The ARUM for a binary treatment of (13) is equivalent to the Heckman-Vytlacil nonparametric selection model for a binary treatment considered, e.g., in Heckman and Vytlacil (1999, 2005), and shown by Vytlacil (2002) to be equivalent to the monotonicity and exogeneity assumptions of Imbens and Angrist (1994). Heckman and Vytlacil (2001) show that the Heckman-Vytlacil nonparametric selection model for a binary treatment results in an identified set for the ATE in (6) and that the model has no identifying power beyond instrument exogeneity for the ATE. Example 4.6 therefore shows that the results of this paper nests the results of Heckman and Vytlacil (2001); Example 4.5, on the other hand, extends their results to nonparametric selection models for a multi-valued treatment. For a partial identification analysis of a class of parameters that includes the ATE under a nonparametric selection model for a binary treatment and sometimes imposing additional restrictions, see, e.g., Mogstad et al. (2018), Han and Yang (2024), and Marx (2024).
@@
Example 4.8.
$Q\{D_{1}=1|D_{0}=1\}$ $=1~{},$ (14)
$Q\{D_{2}=2|D_{0}=2\}$ $=1~{}.$ (15)
$Q\{\mathbbm{1}\{D_{1}=2\}$ $=\mathbbm{1}\{D_{0}=2\}|D_{0}\neq 1,D_{1}\neq 1\}=1~{},$ (16)
$Q\{\mathbbm{1}\{D_{2}=1\}$ $=\mathbbm{1}\{D_{0}=1\}|D_{0}\neq 2,D_{2}\neq 2\}=1~{}.$ (17)
$Q\{D_{0}\neq 0,D_{1}=0\}=0~{}.$
$Q\{D_{0}\neq 0,D_{2}=0\}=0~{},$
$Q\{D_{1}\neq 1,D_{0}=1\}=0~{}.$
It therefore remains to argue by contradiction that
$Q\{D_{1}\neq 1,D_{2}=1\}=0~{}.$ (18)
@@
$=Q\{Y_{1}=y,D_{1}=1|Z=1\}+Q\{Y_{1}=y,D_{0}=1|Z=0\}$
$=P\{Y=y,D=1|Z=1\}+P\{Y=y,D=1|Z=0\}~{},$
where the first two equalities follows from $Q\{D_{0}=D_{1}\}=0$, the third equality follows from Assumption 2.1, and the final equality follows from $Q\in\mathbf{Q}_{0}(P,\mathbf{Q})$. A similar argument establishes identification of $Q\{Y_{0}=y\}$. In contrast, we show in Appendix B.2 that there exists a $P$ for which $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$ and (5) is not a singleton. The identified set for $\theta(Q)$ is therefore not given by (5). We further show in Appendix B.2 that $\Theta_{0}(P,\mathbf{Q}_{1}^{\ast})$ is not a singleton for the same $P$. Thus, the model $\mathbf{Q}$ does have identifying power beyond instrument exogeneity for $\theta(Q)$.

### v3

@@ 3.1 Identified Sets for $\mathbf{Q}\subseteq\mathbf{Q}_{E,M}^{\ast}$
Suppose $\mathbf{Q}\subseteq\mathbf{Q}_{E,M}^{\ast}$ and $\mathbf{Q}|_{Y}$ satisfies Assumption 3.1. Then, for any $P$ such that $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$,
$\Theta_{0}(P,\mathbf{Q})=\prod_{d\in\mathcal{D}}\left[\beta_{d|z^{\ast}(d)}+y^{L}(1-\sum_{y\in\mathcal{Y}}p_{yd|z^{\ast}(d)})),~{}\beta_{d|z^{\ast}(d)}+y^{U}(1-\sum_{y\in\mathcal{Y}}p_{yd|z^{\ast}(d)}))\right]~{}.$ (5)
We now describe some intuition for Theorem 3.2. Note that the distribution of the data, $P$, only contains information on the distribution of $Y_{d}$ for those individuals who would take treatment $d$ for some value of the instrument. On the other hand, it contains no information on the distribution of $Y_{d}$ for those individuals who would not take that treatment for any value of the instrument. Assumption 2.2 implies that individuals would take treatment $d$ at some value of the instrument if and only if $D_{z^{\ast}(d)}=d$, i.e., when maximally encouraged to do so. Assumption 2.1 implies $\beta_{d|z^{\ast}(d)}=\mathbb{E}_{Q}[\mathbbm{1}\{D_{z^{\ast}(d)}=d\}Y_{d}]$, and thus captures all the information from $P$ relevant to $\mathbb{E}_{Q}[Y_{d}]$, which is the first part of the lower and upper bounds in (5). In contrast, Assumptions 2.1 and 2.2 imply that the probability that an individual would not take treatment $d$ for any value of $Z$ is identified from $P$ to be $Q\{D_{z^{\ast}(d)}\neq d\}=1-\sum_{y\in\mathcal{Y}}p_{yd|z^{\ast}(d)},$ but $P$ contains no information on the distribution of $Y_{d}$ for such individuals. Furthermore, that $\mathbf{Q}|_{Y}$ satisfies Assumption 3.1 implies that the model does not restrict the distribution of $Y_{d}$ for such individuals beyond $y\in\mathcal{Y}$, so that we can set $Y_{d}$ to be any value between $y^{L}$ and $y^{U}$ for these individuals, which constitutes the second part of the upper and lower bounds in (5).
@@
Under the instrument exogeneity and monotonicity assumptions of Imbens and Angrist (1994), Balke and Pearl (1993, 1997) found the same form of the identified set for $\theta(Q)$ as (5) when $\mathcal{Y}=\mathcal{D}=\mathcal{Z}=\{0,1\}$. Theorem 3.2 therefore generalizes the result of Balke and Pearl (1993, 1997) to more than two treatment arms and instrument values, to outcomes taking more than two values, and, more surprisingly, to show that the same identified set holds when imposing possibly stronger restrictions on potential treatments than generalized monotonicity.
@@
$\biggl{[}(\beta_{j|z^{\ast}(j)}-\beta_{k|z^{\ast}(k)})+(y^{L}-y^{U})+y^{U}\sum_{y\in\mathcal{Y}}p_{yk|z^{\ast}(k)}-y^{L}\sum_{y\in\mathcal{Y}}p_{yj|z^{\ast}(j)}~{},\\ ~{}(\beta_{j|z^{\ast}(j)}-\beta_{k|z^{\ast}(k)})+(y^{U}-y^{L})+y^{L}\sum_{y\in\mathcal{Y}}p_{yk|z^{\ast}(k)}-y^{U}\sum_{y\in\mathcal{Y}}p_{yj|z^{\ast}(j)}\biggr{]}~{}.$ (6)
@@
Therefore, when imposing generalized monotonicity, as well as when imposing any restriction implying generalized monotonicity, the ATE of $j$ versus $k$ is only identified “at infinity” (Heckman, 1990; Andrews and Schafgans, 1998) in the sense that identification requires
$\min_{z\in\mathcal{Z}}P\{D\neq j\mid Z=z\}=\min_{z\in\mathcal{Z}}P\{D\neq k\mid Z=z\}=0~{}.$ (7)
In other words, identification of the the ATE of $j$ versus $k$ under generalized monotonicity or under any restriction implying generalized monotonicity requires that there is some value of the instrument such that everyone takes treatment $j$ at that value of the instrument, and some value of the instrument such that everyone takes treatment $k$ at that value of the instrument. In contrast, by imposing restrictions on potential treatments that imply that generalized monotonicity is violated, the ATE can sometimes be identified without (7) even when potential outcomes are unrestricted; see Example 5.1 in Section 5 below.
@@ 3.2 Identifying Power of Generalized Monotonicity
Theorem 3.2 above establishes that, for possibly multi-valued $Y$, $D$, and $Z$, and any model $\mathbf{Q}\subseteq\mathbf{Q}_{E,M}^{\ast}$ such that $\mathbf{Q}|_{Y}$ equals the entire set of distributions on $\mathcal{Y}^{|\mathcal{D}|}$, if $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$, then $\Theta_{0}(P,\mathbf{Q})$ equals (5). We now show that $\Theta_{0}(P,\mathbf{Q}_{E,M}^{\ast})=\Theta_{0}(P,\mathbf{Q}_{E}^{\ast})$ as long as $\mathbf{Q}_{0}(P,\mathbf{Q}_{E,M}^{\ast})\neq\emptyset$, so that the identified set for $\theta(Q)$ assuming instrument exogeneity and generalized monotonicity coincides with the identified set assuming instrument exogeneity alone, as long as both assumptions are consistent with the distribution of the observed data. Our result therefore generalizes Balke and Pearl (1993, 1997), which study the case of binary $Y$, $Z$, and $D$.
@@
[…]
@@
[…]
The following lemma, which relies on the observation in Lemma 2.1, establishes the equivalence between the identified sets in (5) and (8) when $\mathbf{Q}_{0}(P,\mathbf{Q}_{E,M}^{\ast})\neq\emptyset$.
@@
Suppose $\mathbf{Q}_{0}(P,\mathbf{Q}_{E,M}^{\ast})\neq\emptyset$. Then, the sets in (5) and (8) coincide.
@@
For any $P$ such that $\mathbf{Q}_{0}(P,\mathbf{Q}_{E,M}^{\ast})\neq\emptyset$, $\Theta_{0}(P,\mathbf{Q}_{E,M}^{\ast})=\Theta_{0}(P,\mathbf{Q}_{E}^{\ast})$.
Remark 3.3.
An implication of Theorem 3.3 is that the identified set on $\theta(Q)$ under Assumption 2.1 alone will be (5) whether or not one imposes Assumption 2.2 as long as the distribution of the data is consistent with Assumptions 2.1 and 2.2.
@@
We now consider some restrictions on potential treatments that have been considered previously in the literature. In each case, we show $\mathbf{Q}\subseteq\mathbf{Q}_{E,M}^{\ast}$; in particular, these restrictions satisfy generalized monotonicity. We emphasize that frequently $\mathbf{Q}\subsetneqq\mathbf{Q}_{E,M}^{\ast}$. In what follows, it is implicitly understood that Assumption 3.1 is satisfied, so that the model imposes no restriction on potential outcomes. Thus, in each example Theorem 3.1 applies and such restrictions do not provide any identifying power for average potential outcomes or ATEs. In Appendix C, we consider three additional examples: an RCT with a “close substitute” as considered in Kline and Walters (2016); the monotonicity and “irrelevance” assumptions considered in Kirkeboen et al. (2016); and an additive random utility model for a binary treatment.
@@
Consider a multi-arm randomized controlled trial (RCT) with noncompliance, where $Z=d$ denotes random assignment to treatment $d$, $D_{d}=d$ denotes that the subject would comply with assignment if assigned to treatment $d$, and $|\mathcal{D}|=|\mathcal{Z}|$. More generally, not necessarily in the context of an RCT, one can interpret $Z=d$ as encouragement to treatment $d$ and interpret $D_{d}=d$ as the subject would take treatment $d$ if encouraged to do so. In this example, $Q$ satisfies Assumption 2.1 because $Z$ is randomly assigned. We may generalize the “no-defier” restriction of Angrist et al. (1996) as: for each $d\in\mathcal{D}$,
@@
i.e., there is zero probability that a subject would not take treatment $d$ if assigned to (encouraged to take) $d$ but would take $d$ if assigned (encouraged) to some other treatment $d^{\prime}\neq d$. If $Q$ satisfies this generalized no-defier restriction, then Assumption 2.2 holds with $z^{\ast}(d)=d$ for all $d$. This no-defier restriction in particular holds in the context of an RCT with “one-sided non-compliance,” where we assume
@@
Remark 4.1.
As explained in Example 4.1, in a multi-arm RCT with one-sided noncompliance, $P$ will necessarily be consistent with Assumptions 2.1 and 2.2, i.e., $\mathbf{Q}(P,\mathbf{Q}_{E,M})\neq\emptyset$. Following Remark 3.3, Theorem 3.3 therefore implies that the identified set for $\theta(Q)$ under Assumption 2.1 alone will be (5) for such a multi-arm RCT. A further implication is that the identified set under Assumption 2.1 on $\mathbb{E}[Y_{j}-Y_{k}]$ for a multi-arm RCT with one-sided noncompliance depends only on the treatment arms for random assignment to treatments $j$ and $k$.
@@
Assumption 2.2 holds for any $Q$ that satisfies (9). To see this, note that Assumption 2.2 can be expressed as the requirement that for each $d\in\mathcal{D}$, there exists $z^{\ast}(d)\in\mathcal{Z}$ such that
$Q\{\mathbbm{1}\{D_{z^{\ast}(d)}=d\}\geq\mathbbm{1}\{D_{z}=d\}\}=1\text{ for all }z\in\mathcal{Z}~{},$
which is immediately implied by (9). Note, however, that $\mathcal{Z}^{\ast}(d)$ may not be a singleton unless some inequalities in (9) are strict.
@@
Although unordered monotonicity implies Assumption 2.2, the converse is generally false. For example, suppose $\mathcal{Z}=\{0,1,2,3\}$ and $\mathcal{D}=\{0,1\}$. Suppose $\mathbbm{1}\{D_{3}=1\}\geq\mathbbm{1}\{D_{z}=1\}$ w.p.1 under $Q$ for $z\neq 3$ and $\mathbbm{1}\{D_{0}=0\}\geq\mathbbm{1}\{D_{z}=0\}$ w.p.1 under $Q$ for $z\neq 0$, but $Q\{D_{1}=1,D_{2}=0\}$ and $Q\{D_{1}=0,D_{2}=1\}$ are both strictly positive. Then Assumption 2.2 holds with $z^{\ast}(0)=0$ and $z^{\ast}(1)=3$, but unordered monotonicity fails. In particular, $\mathbbm{1}\{D_{1}=1\}$ and $\mathbbm{1}\{D_{2}=1\}$ are not ordered, thus violating (9). We thus conclude that if $\mathbf{Q}$ is defined as the set of distributions that satisfy instrument exogeneity and unordered monotonicity, then $\mathbf{Q}\subsetneqq\mathbf{Q}_{E,M}^{\ast}$.
@@
for $g:\mathcal{Z}\times\mathcal{D}\to\Re$ where $\Re$ is the set of real numbers and a random vector $(U_{d}:d\in\mathcal{D})$, whose distribution is absolutely continuous with respect to the Lebesgue measure on $\Re^{|\mathcal{D}|}$ and $Z\perp\!\!\!\perp((U_{d}:d\in\mathcal{D}),(Y_{d}:d\in\mathcal{D}))$. Hence, $Q$ satisfies Assumption 2.1 by construction. Let $\mathbf{Q}$ denote the set of distributions that are consistent with $(D_{z}:z\in\mathcal{Z})$ being determined by (10) for some $g$ and $(U_{d}:d\in\mathcal{D})$ satisfying these requirements. The model $\mathbf{Q}$ is called an additive random utility model (ARUM). A sufficient condition for $Q\in\mathbf{Q}$ to satisfy Assumption 2.2 is that for each $d\in\mathcal{D}$ there exists $z^{\ast}(d)\in\mathcal{Z}$ such that
$g(z^{\ast}(d),d)-g(z^{\ast}(d),d^{\prime})>g(z,d)-g(z,d^{\prime})\text{ for all }d^{\prime}\neq d\text{ and }z\neq z^{\ast}(d)~{}.$ (11)
We refer to the requirement in (11) as uniform targeting of treatment $d$. The terminology is intended to reflect that there is a value of the instrument that maximizes the gains (in terms of $g$) of choosing treatment $d$ versus any other treatment $d^{\prime}$ uniformly across these other possible values of the treatment. In this sense, that value of the instrument targets treatment $d$ uniformly. We now argue by contradiction that (11) implies (3); hence, if (11) holds for all $d\in\mathcal{D}$, then $Q$ satisfies Assumption 2.2. To this end, suppose that, with positive probability, $D_{z^{\ast}(d)}=d^{\prime}\neq d$ but $D_{z^{\prime}}=d$ for $z^{\prime}\neq z^{\ast}(d)$. Then,
$g(z^{\ast}(d),d^{\prime})+U_{d^{\prime}}$ $\geq g(z^{\ast}(d),d)+U_{d}~{},$
@@
$g(z^{\ast}(d),d)-g(z^{\ast}(d),d^{\prime})\leq g(z^{\prime},d)-g(z^{\prime},d^{\prime})~{},$
which violates (11). A particular example when $|\mathcal{Z}|\geq|\mathcal{D}|$ that satisfies (11) with $z^{\ast}(d)=d$ after a suitable relabelling is
@@
The ARUM for a binary treatment is equivalent to the Heckman-Vytlacil nonparametric selection model for a binary treatment considered, e.g., in Heckman and Vytlacil (1999, 2005), and shown by Vytlacil (2002) to be equivalent to the monotonicity and exogeneity assumptions of Imbens and Angrist (1994). Heckman and Vytlacil (2001) show that the Heckman-Vytlacil nonparametric selection model for a binary treatment results in an identified set for the ATE of the form in (6) and that the model has no identifying power beyond instrument exogeneity for the ATE. Example C.1 in Appendix C shows that an ARUM for a binary treatment will satisfy Assumption 2.2 and therefore the results of this paper nest the results of Heckman and Vytlacil (2001). Example 4.4, on the other hand, extends their results to nonparametric selection models for a multi-valued treatment. For a partial identification analysis of a class of parameters that includes the ATE under a nonparametric selection model for a binary treatment and sometimes imposing additional restrictions, see, e.g., Mogstad et al. (2018), Han and Yang (2024), and Marx (2024).
@@
$=Q\{Y_{1}=y,D_{1}=1\mid Z=1\}+Q\{Y_{1}=y,D_{0}=1\mid Z=0\}$
$=P\{Y=y,D=1\mid Z=1\}+P\{Y=y,D=1\mid Z=0\}~{},$
where the first two equalities follows from $Q\{D_{0}=D_{1}\}=0$, the third equality follows from Assumption 2.1, and the final equality follows from $Q\in\mathbf{Q}_{0}(P,\mathbf{Q})$. A similar argument establishes identification of $Q\{Y_{0}=y\}$. In contrast, we show in Appendix B.2 that there exists a $P$ for which $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$ and (5) is not a singleton. The identified set for $\theta(Q)$ is therefore not given by (5). We further show in Appendix B.2 that $\Theta_{0}(P,\mathbf{Q}_{E}^{\ast})$ is not a singleton for the same $P$. Thus, the model $\mathbf{Q}$ does have identifying power beyond instrument exogeneity for $\theta(Q)$.